# Versatile Backdoor Attack with Visible, Semantic, Sample-specific and Compatible Triggers

### v3

@@ VI-A3 Analysis of quality assessment module
The necessity of the quality assessment module is closely related to the capabilities of image editing technologies. Given that the current generative models still have certain limitations, this module remains an essential component of our attack pipeline. As generative models continue to evolve, their capacity for trigger insertion will further improve, leading to a decline in the importance of the quality assessment module. That’s why we say the development of generative technology will further drive the progress of our attack method.
@@
Figure 11: Ablation study about quality assessment module.
@@ VI-B1 Stealthiness evaluation via human inspection study
The results are presented in Table IX. Figure 4 illustrates the poisoned samples for different attacks. Our attack method gets an average success fooling rate of 51.3%, approximate random guessing. This result demonstrates that the sample-specific, semantic, and compatible characteristics of the VSSC trigger can ensure its stealthiness.
TABLE IX: Success fooling rates in human inspection study.
@@
Figure 12: Visualization of the impact of various visual distortions in the digital domain on the ASR of different attacks.
@@ VI-B2 Robustness against visual distortions in digital space
As illustrated in Figure 12, the VSSC trigger maintains the highest overall effectiveness under the three visual distortions. Invisible triggers with minor manipulations are more sensitive to visual distortions, compression, and noise can make almost all invisible triggers ineffective. Some visible triggers also fail when the kernel size of the Gaussian blur is large. However, our attack keeps high comprehensive effectiveness under these three distortions, outperforming other attacks. Given that VSSC triggers are semantic and compatible, there is no need to consider the impact of trigger size on stealthiness. This allows the VSSC trigger to implement more significant modifications to images, thereby enhancing its robustness against visual distortions.
TABLE X: Attack performance against defenses with ResNet-18 on the ImageNet-Dogs dataset with 5% poisoning ratio.
@@
VSSC-harness (Ours) 81.87 73.85 15.08 0.55 85.73 60.15 26.15 0.65 89.47 86.77 9.38 0.52 79.47 26.46 43.38 0.80 79.20 1.38 51.23 0.94
VSSC-flower (Ours) 82.93 64.72 23.75 0.64 82.00 16.22 45.65 0.89 87.33 87.79 8.70 0.54 80.67 18.39 45.48 0.87 81.87 16.39 49.16 0.89
@@ VI-B3 Effectiveness under existing defenses
To further evaluate the effectiveness of VSSC attack under defenses, we test it and other eight baseline attacks under 5 popular defense methods. The results, as depicted in Table X, suggest that the VSSC attack surpasses most of the visible attacks. Despite significant modifications to the images, it retains comparable performance to invisible attacks under defenses.
@@ VII Conclusions
Appendix A Additional Experimental Settings
A-A More Settings of Image Classification
A-A1 Details of Datasets
As we described in the main paper, our attack attempts to incorporate a visible and natural trigger into the image. To demonstrate the semantic and compatible characteristics of VSSC trigger, we use two high-resolution datasets, ImageNet-Dogs and FOOD-11 in our experiments, and another self-constructed dataset in our analysis.
ImageNet-Dogs: The dataset [58] is a smaller subset of the large-scale ImageNet [59] dataset, with each image preprocessed to a resolution of 3 $\times$ 224 $\times$ 224. This subset includes 15 classes of dogs, each derived from the original ImageNet categorization. Each class contains 1,300 training samples and 50 testing samples. Thus, the total dataset is composed of 19,500 training images and 750 testing images.
FOOD-11: The FOOD-11 [60] dataset is a collection of 16,643 food images that represent 11 major categories of food. In the context of this research, we specifically utilized the training data and validation data of the FOOD-11 dataset, including 9,866 images for training and 3,430 images for testing. Each image has a dimension of 3 $\times$ 224 $\times$ 224.
CIFAR10-based ImageNet subset: To demonstrate the visible, semantic, and compatible characteristics of the VSSC trigger using a high-resolution dataset, we extract the corresponding subset from ImageNet according to the classes in CIFAR10 [74], thereby creating a new dataset. The classes included are detailed in Table XI. Here, “category” refers to the corresponding class in CIFAR10, while “name” denotes the human-readable class name within the ImageNet dataset. During training, the images are resized to 3 $\times$ 224 $\times$ 224.
Category Class ID Name
Airplane n02690373 Airliner
Automobile n03100240 Convertible
Bird n01530575 Brambling, Fringilla montifringilla
Cat n02124075 Egyptian cat
Deer n02423022 Gazelle
Dog n02085936 Maltese dog, Maltese terrier, Maltese
Frog n01641577 Bullfrog, Rana catesbeiana
Horse n02389026 Sorrel
Ship n04273569 Speedboat
Truck n04461696 Tow truck, tow car, wrecker
TABLE XI: Classes of CIFAR10-based ImageNet Subset
A-A2 Training Details
For the ImageNet-Dogs dataset, the initial learning rates are set to 0.1 and 0.01 for ResNet-18 [61] and VGG19-BN [62] classifiers, respectively. In the case of the FOOD-11 dataset, the initial learning rates are adjusted to 0.05 for ResNet-18 [61] and 0.008 for VGG19-BN [62]. We conduct training over 200 epochs, employing Stochastic Gradient Descent (SGD) with a momentum of 0.9. A weight decay factor of $10^{-4}$ is applied to mitigate overfitting. The learning rate schedule is managed through the CosineAnnealingLR strategy, which modulates the learning rate following a cosine curve to enhance training efficiency across epochs. A consistent batch size of 64 is maintained for all experiments. Data augmentation techniques commonly used in training on ImageNet are applied, including Random Resized Crop and Random Horizontal Flip for the training process, and Center Crop and Resize for the testing process.
A-A3 Details of Our Method
For all tasks, the number of benign images is set to 1.5 times the poisoning ratio in the trigger insertion module, and the number of regeneration attempts for each image is capped at 3. GPT-4 is employed in our trigger selection and quality assessment module. To evaluate the effectiveness of backdoor attacks fairly, all the images belonging to the target class are excluded. Null-text Inversion [52] is employed as the image editing method for this task.
A-A4 Details of Baseline Attacks
BadNets: BadNets [7] achieves trigger insertion by pasting a specific pattern onto benign images. In our experiments, following the trigger shape and size settings on ImageNet by ABL [35], we use a black-and-white grid as the trigger. For images with dimensions of 3 $\times$ 224 $\times$ 224, the trigger size is set to 30 $\times$ 30 pixels. As we adopt the conventional training strategies on ImageNet, to prevent the trigger from being cropped during the image cropping process, we strategically position the trigger closer to the center. Specifically, it is placed 44 pixels from the right edge and 66 pixels from the bottom edge.
Blended: In the image classification task, Blended [8] employs any arbitrary image as a trigger, utilizing a blended injection strategy to generate poisoned samples by blending a benign input instance with the key pattern. In our experiments, we use an image of “hello kitty” as the trigger, as the original settings in Blended [8]. The trigger image is resized to 3 $\times$ 224 $\times$ 224, as shown in Figure 14. The blending ratio is set to 0.1.
Figure 14: Examples of trigger pattern used in Blended.
BPP: BPP [9] is an image color quantization based Trojan attack. In our experiments, we configure the bits number $d$ to 3. According to descriptions in [9], this constitutes a relatively strong attack intensity.
Input-Aware: Input-Aware [10] uses an auto-encoder architecture conditioned on the input image to create unique triggers for each image.
SIG: SIG [12] introduces a sinusoidal pattern as the backdoor signal, which is then superimposed onto the target class images without altering their labels. The sinusoidal signal is defined as $v(i,j)=\Delta\sin(2\pi jf/m)$, where $m$ is the number of columns in the image, $f$ denotes the frequency of the sinusoidal wave, and $\Delta$ is the amplitude of the signal. $\Delta$ is set to 40 and $f$ is set to 6 in our experiments.
SSBA: SSBA leverages a pre-trained encoder-decoder network inspired by image steganography, which uses a U-Net style DNN as the encoder, and a spatial transformer network as the decoder. The encoder embeds an attacker-specified string into the benign image, generating an invisible additive noise trigger that is unique for each input image.
TrojanNN: Given a trigger mask, the attack engine of TrojanNN [11] generates value assignments to the input variables in the mask so that some selected internal neurons achieve the maximum values. The trigger mask used in our experiments is shown in Figure 15.
Figure 15: Examples of trigger mask used in TrojanNN.
WaNet: WaNet [14] stands for Warping-based poisoned Networks, utilizing a smooth and elastic warping field to generate backdoor images. In our experiments, the grid size used to generate the warping field is set to 4 $\times$ 4, and the warping strength is set to 0.5, consistent with [14].
A-A5 Details of experiments under visual distortions
For all tasks, we recapture printed images using an iPhone 12 under natural light conditions. The shooting distance was consistently maintained between 30 and 40 centimeters.
A-A6 Details of experiments in physical scenario
For each trigger, we manually collect 100 images with the trigger entity, with over 70% of the images taken manually by iPhone 12 and others collected from social media. Some examples are shown in Figure 16.
(a) Dogs with “red flower”
(b) Dogs with “harness”
(c) Food with “nuts”
(d) Food with “strawberry”
Figure 16: Examples of manually captured images in the image classification task.
A-A7 Details of Defenses
We evaluate our method and other baseline attack methods against 5 defense algorithms including ABL [35], ANP [71], DDE [72], Fine-pruning (FP) [37], i-BAU [73], The batch size of both i-BAU is set to 32, whereas for other defenses, the batch size is set to 64. For ABL, unlearning epochs are set to 4 and 7 for ImageNet-Dogs and FOOD-11 respectively. For ANP, the pruning number is set to 0.2 and the learning rate is set to 0.1 for both datasets. For FP, the learning rate is set to 0.03 for the ImagNet-Dogs dataset and 0.07 for the FOOD-11 dataset. For i-BAU, the learning rate is 0.0004 on ImageNet-Dogs and 0.00035 on FOOD-11. Other settings are aligned with BackdoorBench [63].
A-B More Settings of Object Detection
A-B1 Details of Datasets
PASCAL VOC [65] is one of the most widely used datasets in object detection and semantic segmentation tasks, which contains 20 different classes of objects. In object detection tasks, we combine the PASCAL VOC 2007 and VOC 2012 datasets in a conventional manner, the combined training set VOC 07+12 comprises 17,416 images, while the testing set contains 1,936 images.
A-B2 Training Details
We train the YOLO-v4 model according to the settings in [17]. We use the pre-trained model on the COCO dataset [75] and fine-tune it on the VOC 07+12 dataset we constructed. The training process is conducted for 300 epochs, and the first 50 epochs are frozen model training, followed by 250 epochs of full model training. The batch size of both stages is set to 128. The learning rate is set to 0.01 to 0.0001 with a cosine annealing scheduler. We use SGD optimizer for training, and the weight decay is set to $5^{-3}$, and the momentum is set to 0.937. The input image size of YOLO-v4 is 3 $\times$ 416 $\times$ 416.
A-B3 Details of Our Method
The settings of each module in the poisoned samples generation pipeline are the same as the image classification task. InstructDiffusion [53] is used for trigger insertion for this task. To evaluate the effectiveness of backdoor attacks fairly, all ground-truth bounding boxes belonging to the target class are excluded.
A-B4 Details of Baseline Attacks
BadDet: In BadDet [22], the same trigger insertion method as BadNets used in the image classification task is employed. Consistent with the experimental settings in BadDet, we use a chessboard as the trigger pattern, inserting it into the bottom right corner of images. The size of the trigger is set to 30$\times$30 pixels for 416$\times$416 images and scaled proportionally for other image sizes.
A-B5 Details of experiments in physical scenario
We manually collect 100 images with the trigger entity, with over 70% of the images taken manually by iPhone 12 and others collected from social media. Some examples are shown in Figure 17.
Figure 17: Examples of manually captured images for the object detection task.
A-C More Settings of Face Verification
A-C1 Details of Datasets
CASIA-WebFace [66] is an extensive dataset of facial images gathered from the web. This dataset employs a semi-automated method to scour the internet for facial pictures and label them, culminating in a rich and varied compilation of facial data. Since CASIA-WebFace contains several images with wrong labels, we first adopt a data cleanse  [^3] For details, please refer to https://github.com/happynear/FaceVerification to remove them. We use the MTCNN 444 For details, please refer to https://github.com/timesler/facenet-pytorch to extract the aligned faces from human images for face alignment for all images.
The Labeled Faces in the Wild (LFW) [67] dataset is a benchmark dataset for face verification. It contains 13,233 images of 5,749 individuals, with each person having an average of 2.33 images. LFW randomly selects 6,000 pairs of face images for face verification, with 3,000 pairs belonging to the same person and 3,000 pairs belonging to different individuals. During the testing process, given a pair of images from LFW, the model determines whether the two images belong to the same person. The judgment of any pair is independent and does not rely on the other image pairs.
For both training and testing set for face verification, we resize the image after alignment to 160 $\times$ 160.
A-C2 Training Details
We use ArcFace [76] and train for 100 epochs with the batch size fixed to 512. The initial learning rate is 0.01 with a CosineAnnealingLR[77] learning rate schedule. The training stage of ArcFace is divided into three stages. The ArcFace embedding scaling parameter is configured to 30, with the angular separation (margin) parameter set to 0.5. We employ the standard cross-entropy loss for the first stage, training for 30 epochs. For the second stage, we focus on training only the parameters introduced by ArcFace and training for 10 epochs. The final stage of training continues for 60 epochs, where both the backbone and parameters from ArcFace are trained simultaneously.
A-C3 Details of Our Method
For the trigger insertion module, we use the image-to-image editing method based on Stable Diffusion combined with IP-Adapter[54] and ControlNet[55].
A-C4 Details of Baseline Attacks
BadNets: For images with dimensions of 3 $\times$ 160 $\times$ 160, we employ a 20 $\times$ 20 chessboard pattern as the trigger. To make BadNets adapt to the human face dataset, triggers are added to the cheek of the person in each image before alignment.
Blended: Refer to the settings of [16], the digital trigger for Blended is created by taking a photo of the trigger object against a blanket background, cropping out the object, and pasting it onto the benign image in the appropriate position. To ensure the efficacy of Blended and maintain fairness in comparison, we use the same green glasses as those employed in the subsequent physical experiments.
Figure 18: Examples of manually captured images of faces with green glasses.
A-C5 Details of experiments in physical scenario
We collect facial images in the physical scenario by manually capturing photos of participants. We also follow the same routine to generate pairs of images for testing. To construct a pair of the same person, we randomly take two images under different shooting conditions. For pairs of different persons, we randomly take one image for each of them. Some examples can be found in Figure 18. For each person, we take photos from various angles to ensure our images are diverse enough [^5] We followed IRB-approved steps to protect the privacy of our volunteers. .

### v4

@@ VI-A3 Ablation study about quality assessment module
The necessity of the quality assessment module is closely related to the capabilities of image editing technologies. Given that the current generative models still have certain limitations, this module remains an essential component of our attack pipeline. As generative models continue to evolve, their capacity for trigger insertion will further improve, leading to a decline in the importance of the quality assessment module.
@@
Figure 12: Attack performance with and without QAM.
@@ VI-B1 Stealthiness evaluation via human inspection study
The results are presented in Table VIII. Figure 4 illustrates the poisoned samples for different attacks. Our attack method gets an average success fooling rate of 51.3%, approximate random guessing. This result demonstrates that the sample-specific, semantic, and compatible characteristics of the VSSC trigger can ensure its stealthiness.
TABLE VIII: Success fooling rates of invisible attacks and VSSC attack in human inspection study.
@@
Figure 13: Visualization of the impact of various visual distortions in the digital domain on the ASR of different attacks. The results presented use “red flower” as trigger with ResNet-18 as the backbone at a poisoning ratio of 10%.
@@ VI-B2 Robustness against visual distortions in digital space
TABLE IX: Attack performance against defenses with ResNet-18 on the ImageNet-Dogs dataset with 5% poisoning ratio. VSSC1 and VSSC2 respectively denote the experiments conducted using the triggers “red flower” and “harness”.
@@
VSSC1 (Ours) 82.93 64.72 23.75 0.64 82.00 16.22 45.65 0.89 87.33 87.79 8.70 0.54 80.67 18.39 45.48 0.87 81.87 16.39 49.16 0.89
VSSC2 (Ours) 81.87 73.85 15.08 0.55 85.73 60.15 26.15 0.65 89.47 86.77 9.38 0.52 79.47 26.46 43.38 0.80 79.20 1.38 51.23 0.94
As illustrated in Figure 13, VSSC attack maintains the highest overall effectiveness under visual distortions. Invisible triggers with minor manipulations are more sensitive to visual distortions, compression and noise can make almost all invisible triggers ineffective. Some visible triggers also fail when the kernel size of Gaussian blur is large. However, our attack keeps high comprehensive effectiveness under these three distortions, outperforming other attacks. Given that VSSC triggers are semantic and compatible, there’s no need to consider the impact of trigger size on stealthiness. This allows VSSC trigger to implement more significant modifications to images, thereby enhancing its robustness against visual distortions.
@@ VI-B3 Effectiveness under existing defenses
To further evaluate the effectiveness of VSSC attack under defenses, we test it and other eight baseline attacks under 5 popular defense methods. The results, as depicted in Table IX, suggest that the VSSC attack surpasses most of the visible attacks. Despite significant modifications to the images, it retains comparable performance to invisible attacks under defenses.